# ERABAL: Enhancing Role-Playing Agents through Boundary-Aware Learning

### v1

@@ 4.4.2 Comprehensive Evaluation Results
5 Further Analysis
5.1 Effect of Model Scale
We further study the performance variation with Erabal on models of different scales. Experiments are conducted on LLaMA-7B, LLaMA-13B, and LLaMA-33B. Figure 4 shows the experimental results, from which we observe that the performance on the four metrics improves as the model size increases. However, for Cons. and Rej., the improvements tend to plateau when the model size exceeds 13B. Regarding role consistency and knowledge tasks, since these two tasks are relatively difficult, the performance improvements are more significant.
5.2 Effect of Data Scale
We also conduct experiments based on LLaMA2-7B with 1/8, 1/4, 1/2, and full-sized datasets to study the impact of data scale. As depicted in Figure 5, as the data volume increased, almost all metrics showed rapid improvement. We believe that constructing more boundary-aware dialogues through Erabal will yield even better results. However, due to resource limitations, we limit our dataset to a maximum of approximately 16k dialogue data in this experiment. The performance on the Know. metric shows more significant improvement compared to the other three metrics, which is consistent with our previous analysis.
@@ 7 Conclusion
In this work, we propose Erabal, a boundary-aware learning framework aimed to enhance the RPLAs’ role-playing capabilities. Through Erabal, we collect an effective role-playing dataset and train RPLAs based on popular generalist models. Evaluation results demonstrate that our method significantly improves the RPLAs’ performances across all metrics. Further analysis reveals that by increasing model parameters and data scale, our method can achieve even better performance.
@@ A.1 Prompts
All prompts are listed in Tables 816. The prompt of Single-Agent can be found in Table 15.
@@ Appendix B Experimental Details
B.1 RoleBench
RoleBench (Wang et al., 2023c) is a role-related dataset designed to enhance the role-specific knowledge density within synthetic instruction datasets. It involves a meticulous three-step process: segmenting role profiles, generating question-confidence-answer triplets, and filtering and post-processing data to ensure high-quality, role-specific instruction data. This dataset is particularly adept at handling the nuances of script-agnostic and script-based instructions, utilizing a confidence score to improve question quality and address potential issues such as incompleteness or hallucinations.
B.2 WikiRoleEval
This WikiRoleEval meticulously designs metrics to assess three critical aspects of role-playing LLMs’ performance: consistent role identity, accurate role-related knowledge, and the ability to recognize and reject unknown questions.
For consistent role identity, WikiRoleEval suggests evaluating an LLM’s ability to maintain a designated character’s consistency across a multi-turn dialogue. This is achieved by structuring the assessment as a multi-choice problem with four potential role candidates, where an additional LLM judger determines the most suitable character based on the dialogue context.
Regarding accurate role-related knowledge, the framework emphasizes the importance of LLMs conveying role-specific knowledge accurately, avoiding factual errors. WikiRoleEval addresses the challenge of factual assessment by using a dialogue-simulation scheme to capture the essential knowledge for each dialogue round, enabling a judging LLM to evaluate the appropriateness of a response in integrating consistent knowledge.
Lastly, for unknown question rejection, WikiRoleEval introduces a metric to evaluate an LLM’s ability to reject questions outside a character’s cognitive edge, enhancing the realism and immersion of the role-play. This involves annotating questions based on the cognitive edge of each character and employing an LLM judger to assess the model’s rejection accuracy.
@@ B.3 CharacterEval
B.4 Training Implementation Details
We include the role-playing instruction in the model’s system instruction, which comprises the role name, role characteristics, and format requirements. In the case of Baichuan2-Chat-13B, the format is {system instruction}<reserved_106>{input}<reserved_107> {response}</s>. Details of the role-playing system instruction format can be found in Table 8.
@@ Appendix D Complete Examples
Knowledge-Hallucination(KH) This metric evaluates the model’s ability to avoid generating responses with fabricated knowledge, maintaining the role’s identity consistency.
@@
Prompt for Role-Consistency Evaluation
@@
Table 17: The prompt for role-consistency evaluation.
@@
Baichuan2-13B-Chat(+Erabal +BPO): Golly, that sure sounds like a hootenanny! But I gotta admit, I prefer my tires spinning on the racetrack over sipping on some bubbly any day.
@@
Baichuan2-13B-Chat(+Erabal +BPO): Ah, such situations are indeed intriguing. One must learn to navigate through the complexities of human emotions and ambitions. As someone who has experienced a turbulent life, I understand the dynamics of such arguments all too well.

### v2

@@ 7 Conclusion
In this work, we propose Erabal, a boundary-aware learning framework aimed to enhance the RPLAs’ role-playing capabilities. Through Erabal, we collect an effective role-playing dataset and train RPLAs based on popular generalist models. The evaluation results demonstrate that our method significantly improves the performance of RPLAs across all metrics, and it remains consistent with human evaluation results. Further research indicates that increasing model parameters and data scale leads to performance improvements, but the effects vary across different metrics.
@@ A.1 Prompts
All prompts are listed in Tables 1119.
@@ Appendix B Experimental Details
B.1 WikiRoleEval
WikiRoleEval meticulously designs metrics to assess three critical aspects of role-playing LLMs’ performance: consistent role identity, accurate role-related knowledge, and the ability to recognize and reject unknown questions. For consistent role identity, WikiRoleEval suggests evaluating an LLM’s ability to maintain a designated character’s consistency across a multi-turn dialogue. This is achieved by structuring the assessment as a multi-choice problem with four potential role candidates, where an additional LLM judger determines the most suitable character based on the dialogue context. Regarding accurate role-related knowledge, the framework emphasizes the importance of LLMs conveying role-specific knowledge accurately, avoiding factual errors. WikiRoleEval addresses the challenge of factual assessment by using a dialogue-simulation scheme to capture the essential knowledge for each dialogue round, enabling a judging LLM to evaluate the appropriateness of a response in integrating consistent knowledge. For unknown question rejection, WikiRoleEval introduces a metric to evaluate an LLM’s ability to reject questions outside a character’s cognitive scope, enhancing the realism and immersion of the role-play. This involves annotating questions based on the cognitive edge of each character and employing an LLM judger to assess the model’s rejection accuracy.
@@ B.2 CharacterEval
B.3 Training Implementation Details
We include the role-playing instruction in the model’s system instruction, which comprises the role name, role characteristics, and format requirements. Details of the role-playing system instruction format can be found in Table 11.
@@ B.3 Training Implementation Details
B.4 Ablation Study
We investigate the contributions of different components in our Erabal framework. The components we analyze include the dialogue planner, topic manager, and information verifier. A conceivable alternative is to employ simpler prompt engineering methods, such as the Chain-of-Thought (COT) method, to achieve similar results without the complexity of workflow. To explore this, we attempt to use the COT method for data construction. However, as indicated by the results in Table 7, the diversity of the generated dataset is significantly lower. In contrast, by incorporating both the dialogue planner and the topic manager, we ensure that the generated dialogues do not remain constrained to the same topics, leading to notable improvements in dialogue quality and diversity. The prompt of Single-Agent can be found in Table 18.
Method BERTScore $\downarrow$ Dist-1 $\uparrow$ Dist-2 $\uparrow$ Dist-3 $\uparrow$
Single Agent (COT) 0.644 0.457 0.829 0.960
Ours 0.591 0.751 0.930 0.969
   w/o Topic Manager 0.768 0.411 0.805 0.953
   w/o Dialogue Planner 0.771 0.440 0.821 0.968
   w/o Information Verifier 0.794 0.403 0.792 0.940
Table 7: Ablation study results. BERTScore and Distinct-n metrics for various configurations of our model. Lower BERTScore and higher Distinct-n in the corpus indicate improved diversity.
B.5 Human Assessments
We develop a human evaluation framework based on WikiRoleEval (Section 3.2.1), referred to as WikiRoleEval-Human, where the LLM evaluator is replaced with human. For this assessment, we randomly select one-fifth of the generated data in both Chinese and English and compare several key models. Table 8 shows the results, which are consistent with those obtained from the automatic evaluation presented in Table 4. However, for the Consistency (Cons.) metric, human judgment surpassed GPT-4’s evaluation, yielding more concentrated scores. This is because the ability of humans to complete role-related multiple choice questions is far superior to LLMs.
Models Cons. Know. Rej.
GPT-4-Turbo 0.84 6.82 0.79
CharacterGLM 0.91 4.24 0.75
Baichuan2-13B-Chat 0.83 4.19 0.67
   +RoleBench 0.87 4.17 0.69
   +Ditto 0.89 4.84 0.74
   +Erabal (SFT) 0.93 5.20 0.79
   +Erabal (SFT+DPO) 0.95 5.34 0.80
Table 8: Human evaluation results using the WikiRoleEval-Human framework, comparing several conversational models. The metrics include Consistency (Cons.), Knowledge (Know.), and Rejection of incorrect information (Rej.).
The human assessment results indicate that models trained using the Erabal framework (SFT and SFT+DPO) outperform the baseline models, particularly in terms of dialogue consistency and rejection of incorrect information. The Erabal (SFT+DPO) model, in particular, achieved the highest scores across all metrics, highlighting the strength of our boundary-aware learning approach in generating coherent and factually accurate responses.
B.6 Direct Usage of GPT-4 with Prompts
In this paper, we propose the Erabal framework based on GPT-4 to generate high-quality training data, which is then used to train downstream models. A natural question arises: why not directly use this framework—GPT-4 with a well-designed prompt—as the model itself, bypassing the need for further training?
Theoretically, it is feasible to use GPT-4 directly with a role-playing prompt. To investigate this, we conduct experiments to evaluate GPT-4’s performance under a carefully designed prompt, using our Boundary Scenario Evaluation method (Section 3.1). The results in Table 9 show that GPT-4 with a well-designed prompt achieves high performance, with a score of $0.6420.642$, compared to the baseline GPT-4 score of $0.5390.539$.
However, while these results appear promising, they are not practical for real-world applications. The improvement in performance can be largely attributed to the fact that the prompt contains explicit hints or reminders, which act as "cheating" mechanisms. For example, the prompt often includes statements alerting GPT-4 that certain queries contain false or misleading information, thereby guiding the model toward the correct response. This type of prompt engineering, while effective in controlled experimental settings, is unrealistic in typical human-machine interactions, where users do not provide such detailed guidance.
Thus, while GPT-4 with a well-designed prompt can perform well, our goal is to create a framework that generates training data to train models capable of robust, independent reasoning in dynamic, real-world environments—without relying on artificial cues embedded in the prompt.
The prompt used in the evaluation is as follows:
Role-playing Prompt
{seed_feature}
Please note that the query contains false information:
{counterfactual_information}
Method Boundary Scenario Evaluation
GPT-4 0.539
GPT-4 with prompt 0.642
Table 9: Boundary Scenario Evaluation results comparing GPT-4 with a standard role-playing prompt and GPT-4 with a well-designed prompt.
@@ Appendix D Complete Examples
Knowledge-Hallucination(KH) This metric evaluates the model’s ability to avoid fabricating knowledge, with a focus on maintaining consistency in world knowledge rather than character-specific knowledge.
@@
Prompt for Role Consistency Evaluation
@@
Table 20: The prompt for role consistency evaluation.
@@
Baichuan2-13B-Chat(+Erabal +DPO): Golly, that sure sounds like a hootenanny! But I gotta admit, I prefer my tires spinning on the racetrack over sipping on some bubbly any day.
@@
Baichuan2-13B-Chat(+Erabal +DPO): Ah, such situations are indeed intriguing. One must learn to navigate through the complexities of human emotions and ambitions. As someone who has experienced a turbulent life, I understand the dynamics of such arguments all too well.